# Mitigating Heterogeneity in Federated Multimodal Learning with Biomedical Vision-Language Pre-training

### v1

@@ 5.4 Analysis experiment
In this section, we mainly conduct analysis experiments using the SLIA backbone method for pre-training unless specifically stated.
@@
Strategy $1\%$ (cls.) Chest imaG. (ret.)
@@
[…]
ours 83.21 80.50 29.49 26.99 73.03 67.65
Strategy O.O.D In domain In domain Wst.
Rec.@1 Rec.@5 Rec.@1 Rec.@5 Rec.@1 Rec.@5
FedX 12.31 61.04 3.22 16.76 1.58 11.94
FedMAE 12.93 63.72 5.05 23.91 4.34 20.86
FedEMA 15.19 66.94 24.01 67.05 21.85 62.44
FedAvg 15.3 66.62 28.83 72.12 25.34 66.68
FedRGB (Ours) 15.43 67.29 29.49 73.03 26.99 67.65
Figure 8: The trade-off adjusted by the hyper-parameter $\alpha$.
Figure 9: The role of the hyper-parameter $\alpha$.
Figure 10: The role of the radius of the uncertainty set $\rho$.
Robust teacher align module is better than simply averaged one. To investigate the efficiency of the robust align module, we replace it with an align module that is simply averaged by local ones and pre-train a new model. As shown in table 4, the downstream performance of the replaced one is lower than the original one. This means the robust align module can enhance the ability to capture cross-modal alignments and learn transferable representation. Strong regularization is better than weak regularization for guiding the updation. To investigate the effect of the guidance from the robust align module, we replace the L-2 regularization with contrastive loss as a weaker regularization and pre-train a new model. As shown in table 4, the weaker guidance leads to lower classification accuracy and retrieval accuracy in the downstream tasks. Guidance hyper-parameter adjusts the average and disparity of client performances. Figure 8 shows that higher guidance achieves smaller disparity while bringing less harm on average performance. We changed the hyperparameter $\alpha$ from 2 to 10, the disparity of the retrieval accuracy on test sets of different clients decreased significantly while there was only a slight loss on the averaged performance. On the other hand, Figure 9 shows that increasing $\alpha$ wouldn’t result in a penalty in the transferability of the pre-trained backbone, the downstream task performances are maintained at a high level as we enlarge the guidance. Detailed results can be seen in Table 15.
Figure 11: Robustness & convergence on the number of com. turns.
Convergence check: The downstream performance of the pre-trained got promoted as we increased the number of comm. turns. After 20 communications, the performance turned out to be stable which verified the convergent ability of our method. Detailed results can be seen in Table 14. Our method has better classification accuracies on unseen out-of-distribution multi-modal X-Ray datasets. The multi-modal retrieval accuracy in the unseen domain of our method beats the baseline by a large margin as shown in Table 5, which verifies our method’s ability to learn generalizable cross-modal alignment. Meanwhile, our FedRGB has higher in-domain averaged and worst-case acc.. Therefore, improvements in the in-domain performance of FedRGB do not come at the cost of decreased out-of-domain performance. Smaller radius of uncertain set brings higher generalization and transferability on downstream task. Figure 10 shows the downstream performance of models pre-trained with different radius of uncertainty set in the DRO process. As the $\rho$ increases, the performance of fine-tuning-based downstream tasks shows an upward trend. Detailed results can be seen in Table 10.
@@ 6 Conclusion
Data limitation is a long-standing problem in the multi-modal learning domain. Despite federated learning can leveraging datasets from multiple sources while guaranteeing privacy issues, its performance would be damaged by data heterogeneity. Inspired by our empirical findings on the impact of heterogeneity on federated multi-modal learning, we propose the FedRGB framework to multigate heterogeneity for federated biomedical vision-language pre-training. The efficiency of our method has been verified by comprehensive experiments.
7 Limitation
The design of the proposed FedRGB approach has taken it into serious consideration to mitigate the potential negative societal impact.. For example, to tackle the data privacy issue, FedRGB has avoided the transmission of any data or latent representations during all server-client communications, while only transferring model parameters and loss, thereby minimizing the risk of privacy leakage in federated learning as much as possible. However, our model may still be subject to privacy-related issues due to the communication of model parameters, which is a common problem faced by other federated learning methods as well. There are also research studies in this line to avoid privacy leakage from model parameters. For fairness, our method aims at optimizing the performance of the worst-case client to promote model equity across different cohorts, whichhas shown great generalizable ability on unseen data. This approach ensures that the pre-trained model not only performs well and fairly on the data from hospitals participating in the federated training but also demonstrates good performance on hospitals that did not participate in the training. Meanwhile we acknowledge that there may still exist limitations. Firstly, If we can resolve the trade-off between privacy protection and data representation communication, it may allow us to design more advanced algorithms to further enhance the performance of federated multimodal pre-training models. Furthermore, in this work we mainly consider the multimodal medical datasets of X-Ray imageeport due to the limitation of the public datasets. However, some rare diseases and more complicated modalities in various medical applications lack paired data, which] may bring other challenges for federated multimodal pretraining.
@@ 8 Acknowledgement
Appendix A Problem Formulation
We aim at mitigating the data heterogeneity issues in federally pre-training biomedical vision-language models to better capture the semantic alignment between the two modalities, with heterogeneous private datasets from different clients. In this section, we will formulate our problem set-up, assumptions, and related notations. Federated Vision-Language Pre-training. In this paper, we consider the federated vision-language pre-training (VLP) problem in the biomedical domain. Suppose we are given $N$ clients $\{C_{i}\}_{i=1}^{N}$, each client has a corresponding local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues. Without access to centralized pre-training, federated pre-training aims to conduct vision-language representation learning across local clients, while still capture a robust and generalizable cross-modal alignment, by efficiently aggregating client models $\{M_{i}\}_{i=1}^{N}$ trained on their private datasets. Similar to conventional federated learning scenarios, our federated pre-training consists of multiple communication turns. For communication turn $t$, client models $\{M_{i}\}_{i=1}^{N}$ completed the same number of local iteration steps respectively, and the server receives and aggregates the weights sent from local models to produce the aggregated server model $M^{t}$. Then it will be sent back to clients and re-trained on $\{D_{i}\}_{i=1}^{N}$ separately to get new local models $\{M_{i}^{t+1}\}_{i=1}^{N}$. The federated vision-language pre-training is a more practical and general task compared to related works. Table 1 lists the most similar settings in previous works, and the difference between our task and theirs. Data Formulation. We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n},Y=\{y_{i}\}_{i=1}^{m},n,m\in N^{+}$. We assume there exists latent variables U that embed semantics across two modalities such that $X=T_{x}(U),Y=T_{y}(U)$, where $T_{x}(\cdot),T_{y}(\cdot)$ are assumed transformations that is not explicit to know. For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report. Thus, we assume each client has paired image-text data that are implicitly correlated through the same latent variables. Therefore, local client datasets could be written as $\mathcal{D}:\{D_{i}:(T_{x}(u),T_{y}(u)),u\sim P_{i}(U)\}_{i=0}^{N}$. Here, we assume the local datasets are heterogeneous, which means the distribution of latent variables $P_{i}(U)$ is different across clients. We formulate two distributions $\mu_{i}\in P(X_{i}),\nu_{i}\in P(Y_{i})$ following (Chen et al., 2020a) as $\mu_{i}=\sum_{j=1}^{n_{i}}a_{i}\delta_{T_{x}(u_{j}))_{i,j}}$ and $\nu_{i}=\sum_{j=1}^{n_{i}}b_{j}\delta_{T_{y}(u_{j})_{i,j}}$, with $\delta_{x}$ as the Dirac function centered on $x$. The weight vectors $a_{i}=\{a_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ and $b_{i}=\{b_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ belong to the $(n-1)$- dim simplex, where $\mu_{i}$ and $\nu_{i}$ are probability distributions. This formulation can also be applied to feature vectors or representations. For a given set of distributions $\{P_{i}\}_{i=1}^{N}$, we can form a new distribution $P^{c}=\sum_{i=1}^{N}w_{i}P_{i},\textbf{w}\in\Delta$ where $\Delta$ is a $(N-1)$-dim simplex by weighted averaging the sample vectors. Model Architectures. Following the VLP works (Radford et al., 2021; Yang et al., 2022), we frame the VLP using a dual-encoder structure, where the model for each modality consists of a feature encoder module $f(\cdot)$ that outputs feature vectors given the raw input, and a feature alignment module $g(\cdot)$ that projects feature vectors to aligned representations in the shared embedding space. Denote the feature vector as $s$, and the aligned representation as $z$. We model the outputs of the encoder and alignment modules as conditional distributions. Specifically, the encoder takes the multi-modal raw input to generate semantic embeddings of the corresponding modality, which could be formulated as $P(s|U)$. Similarly, that of the alignment module can be formulated as $P(z|S,U)$.
Appendix B Details of Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
In order to guide the local client training, we need to provide a teacher alignment module as mentioned in the last section, which can capture globally unbiased and generalizable cross-domain alignment. How can we obtain such a teacher guidance model? We propose to learn the unbiased cross-domain alignment by considering the robustness across various distributions of local clients. A teacher alignment module should avoid distorting the aggregated feature encoder while learning the cross-modal alignments. Additionally, we introduce a much harder task to force the module to be more robust to biased alignment, by considering the influence of the biased alignment module on the feature encoder. Furthermore, to learn a teacher model that performs well on varies data distributions, we minimize distortion degree under the most biased local alignment with the “worst-case” local data distribution. Inspired by the distributionally robust optimization (DRO) (Cotter et al., 2019), we model the de-centralized pre-training framework as a two-stage DRO learning process. In the first stage, the local models are trained with $\mathcal{L}_{local}$ (Eq. 2), while the whole model is jointly optimized with the guidance of the teacher alignment module. In the second stage, we learn a robust teacher alignment module that performs well conditioned on the worst-case local models where the feature encoder is distorted the most severely. We aim to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local aligner and local data distribution, which represents possible cases that the model would encounter. We construct the uncertainty set based on the distribution of client datasets and local aligners learned in the first stage. From the probabilistic view, the local aligners could be written as $P(z|U,f)$, which takes the output of the given encoder $f$ to generate the aligned representations, while $f$ takes the raw input from the local data distribution $P(U)$ to generate the semantic embedding that has not well learned the cross-modal alignment. Therefore, the distribution to be considered should be $P(z,U|f)$. Hence the uncertainty set can be written as:
$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i}(z,U|f),w\in\Delta_{N-1}$ (7)
where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability. Finally, with the teacher align module denoted as $g^{*}$, indicating the aggregated encoder after the first stage training, the learning object in the second stage can be formulated as:
$\footnotesize min_{g^{*}}\ max_{P\in Q^{c}}\ \mathcal{L}_{dro}(g^{*},P,f^{*})$ (8)
Following (Zhang et al., 2023b), the optimization goal can be constructed by distributions $P_{i}$ of the support set:
$\footnotesize min_{g^{*}}\ max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (9)
where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ are the losses calculated on the local clients respectively. While we learn $g^{*}$ by gradient descent, the mirror gradient descent process of optimizing w could be solved efficiently (Zhang et al., 2023b; Duchi et al., 2008):
$\footnotesize{w_{i}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}e^{\gamma v_{i}}}/\sum_{i=1}^{N}{w_{i}e^{\gamma v_{i}}},\rho\right)$ (10)
where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$. We alternatively optimize the w and the teacher align module. For each communication turn of the federated learning, we first obtain an averaged encoder, and a set of local align modules in the first stage. Then we solve the weights and optimize the parameters of the teacher align module. Hence the w could be written in a more trackable form as:
$\footnotesize{w_{i}^{(r+1)}}=\text{Proj}{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (11)
where $r$ refers to the communication turn.. Finally, we formulate $\mathcal{L}_{dro}$ to measure the performance of the teacher align module of alleviating distortion on the encoder. Based on equations (3) and (4), we update the teacher alignment module in local clients with weights w. Specifically, given a client $c$, we use $f^{*}$ learned in the first stage of the current communication turn to generate features $(\tilde{v^{c}},\tilde{l^{c}})$, and a learnable copy of $f^{*}$ to generate distorted features $(v^{c},l^{c})$. The distorted features are input to the frozen $g$ to calculate the task loss $L_{task}$, and are input to the $g^{*}$ to get calibrated representations $(z_{I}^{c*},z_{T}^{c*})$. Meanwhile, $(\tilde{v^{c}},\tilde{l^{c}})$ are input to $g^{*}$ to get representation $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$ for regularization. The regularization term could be written as:
$\footnotesize\mathcal{L}_{dro,\tau}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z^{c*}}_{I,i}-z^{c*}_{I,i})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z^{c*}}_{T,i}-z^{c*}_{T,i})^{2})+\mathcal{L}_{CL,\tau}$ (12)
where $\mathcal{L}_{CL}$ is the contrastive loss (Radford et al., 2021) between $(z^{fz,c*}_{img,i},z^{fz,c*}_{txt,i})_{i=0}^{N_{c}}$, and $N_{c}$ refers to the number of samples in the client $c$.
Appendix C Experiment Details
C.1 Pre-training setup
@@
Following (Wang et al., 2022) we utilize the MIMIC-CXR (Bigolin Lanfredi et al., 2022) dataset for multi-modal pre-training. This dataset is widely used in the biomedical multi-modal learning domain, with $227,835$ image-text pairs from $65,379$ patients. Some related works also have imported additional features to image-text pairs to augment the data. However, we only use the image-text pairs for pre-training to make the results more generalizable. During the pre-training, local clients only have access to their highly heterogeneous datasets. To construct the heterogeneous client datasets, following (Yan et al., 2023) we employ the Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to divide the MIMIC-CXR dataset into $5$ partitions based on a selected sensitive attribute. For implementation, we import the corresponding attribute information of given image-text pairs from the Chest-Imagenone dataset (Wu et al., 2021) and divide local datasets based on disease category. The disease category is a multi-label binary attribute and is transformed into a multi-class label. That’s because the words in the clinical report are highly related to the disease category as illustrated in 12. We set the heterogeneity degree in the LDA algorithm to be 1 and the subgroups of the MIMIC-CXR that are used as local datasets are summarized in Table 7. Specifically, we select 5 commonly considered diseases (Bannur et al., 2023): ’Edema’,’Pleural Effusion’, ’Consolidation’, ’Pneumothorax’, and ’Pneumonia’. We set the non-NaN value to 1 and then set NaN value to 0 to construct a 5-way binary multi-label. Then we get $2^{5}$-category multi-class label, and we run LDA on it. The distribution of the disease of the client dataset is summarized in Table 7.
Figure 12: Illustration of the divided heterogeneous subgroups for each client.
FedAvg Test on C1 Test on C2 Test on C3 Test on C4 Test on C5
Retrain On C1 0.092 0.078 0.08 0.073 0.082
Retrain On C2 0.054 0.049 0.046 0.047 0.048
Retrain On C3 0.069 0.072 0.07 0.066 0.067
Retrain On C4 0.092 0.09 0.084 0.093 0.088
Retrain On C5 0.073 0.08 0.07 0.068 0.087
Avg 0.076 0.0738 0.07 0.0694 0.0744
@@
We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 client datasets. Each divided dataset consists of train splits and validation splits based on the notation of the MIMIC-CXR. We use $404\times A$ 40 for pre-training and set the batch size to 388. We set the learning rate to $2\times 10^{-5}$, the number of communications to 25. We set the updating step size $\gamma=1$, and uncertainty radius $\rho=1$, hyper-parameter $\alpha=1,\beta=2$ in the main experiments. For each communication, we randomly sample 50 batches of data from the client datasets. To obtain the guidance, we run an extra communication for our method.
C.2 Downstream tasks
We conduct Few-shot classification, Medical image segmentation, and Image Retrieval downstream tasks to evaluate the generalization ability of the pre-trained model. Few-shot classification. To verify the efficiency of the pre-trained model on the general medical image task, we test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) (Shih et al., 2019), the task is to predict if the image reflects a pneumonia-positive phenomenon. (2) Covidx (Wang et al., 2020), which has three categories that COVID-19, non-COVID pneumonia and normal. We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset of downstream tasks, and evaluate the image classification accuracy. For evaluation, we report the classification accuracy of the models on them. Medical image segmentation. To investigate the transfer ability of our model on fine-grained tasks, we conduct experiments on medical image segmentation on the RSNA (Wang et al., 2020) benchmark. We follow (Wang et al., 2022) to convert object detection ground truths of RSNA into masks for semantic segmentation. Similar to (Huang et al., 2021), we construct a U-Net framework using a decoder and our pre-trained image encoder. During the fine-tuning, we freeze the encoder and fine-tune the decoder using 1%, 10% of the training data of the downstream dataset, and then use the Dice score to evaluate the performance of our model. Image retrieval. To verify if the pre-trained models have learned the semantics and alignment relationships between the image and text contained in the pre-training dataset, we conduct the image retrieval downstream task. Specifically, we utilize the validation splits of the local clients to test the image retrieval performance of the aggregated models. For a text from a given batch of input image-text pairs, we calculate the similarities between this text and images in this batch. Then we sort these similarities and select the top-1 and top-5 images from them. If the corresponding image of the text is contained in the selected set, then the image is retrieved correctly. Here we utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models. Out-of-Distribution testing benchmark. We adapt the EHRXQA (Bae et al., 2023) VQA dataset for evaluation. This dataset is built on the meta-data notation of MIMIC-CXR (Bigolin Lanfredi et al., 2022) and ChestImaGenome (Wu et al., 2021). We merge the question and answer sentences as the text input and use the corresponding image to construct the image-text pairs for the multi-modal retrieval task. Other settings are the same as the ordinary image retrieval task.
C.3 Backbones and Baselines

### v2

@@ 5.4 Analysis experiment
In this section, we conduct analysis experiments using ConVIRT as backbone pre-training method.
@@
Alignment Module Guidance $1\%$ (cls.) Chest imaG. (ret.)
@@
Robust teacher align module is better than simply averaged one. To investigate the efficiency of the robust align module, we replace it with an align module that is simply averaged by local ones and pre-train a new model. As shown in table 3, the downstream performance of the replaced one is lower than the original one. This means the robust align module can enhance the ability to capture cross-modal alignments and learn transferable representation.
Strong regularization is better than weak regularization for guiding the updation. To investigate the effect of the guidance from the robust align module, we replace the L-2 regularization with contrastive loss as a weaker regularization and pre-train a new model. As shown in table 3, the weaker guidance leads to lower classification accuracy and retrieval accuracy in downstream tasks.
Guidance hyper-parameter adjusts the average and disparity of client performances. Fig. 3.a shows that higher guidance achieves smaller disparity while bringing less harm on average performance. We changed the hyperparameter $\alpha$ from 2 to 10, the disparity of the retrieval accuracy on test sets of different clients decreased significantly while there was only a slight loss on the averaged performance. On the other hand, Fig. 3.b shows that increasing $\alpha$ wouldn’t result in a penalty in the transferability of the pre-trained backbone, the downstream task performances are maintained at a high level as we enlarge the guidance. Detailed results can be seen in Table 13.
Smaller radius of uncertain set brings higher generalization and transferability on downstream task. Fig. 3.c shows the downstream performance of models pre-trained with different radius of uncertainty set in the DRO process. As the $\rho$ increases, the performance of fine-tuning-based downstream tasks shows an upward trend. Detailed results can be seen in Table 7.
Convergence. We show the transferability of our models pre-trained with different communication turns in Fig. 3.d. As commun. turn reaches 20, the downstream performances tend to converge.
We have also conducted experiments to explore other properties of our proposed method. We have found that (1). Our method has successfully avoided the distortion caused by heterogeneous datasets during the local training. (2). Bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model. (3). Our method is more generalizable on out-of-distribution multi-modal X-Ray datasets. Detailed results can be seen in Appendix C
@@ 6 Conclusion
Data limitation is a long-standing problem in the multi-modal learning domain. Despite federated learning can leveraging datasets from multiple sources while guaranteeing privacy issues, its performance would be damaged by data heterogeneity. Inspired by our empirical findings on the impact of heterogeneity on federated multi-modal learning, we propose the FedAID framework to multigate heterogeneity for federated medical vision-language pre-training. The efficiency of our method has been verified by comprehensive experiments. While introducing representation from other clients might bring larger improvement, we still consider the most privacy-preserved setting where representations are not transmissible. Further work could explore how to introduce data diversity while keeping local data private.
@@ 7 Acknowledgement
Appendix A Empirical Finding
In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our FedAID. We conduct experiments on the image-text retrieval task, which reflects the ability to capture vision-language alignments. We utilize the ConVIRT backbone strategy and consider the performance of the centralized pre-trained model (Global), the server model pre-trained in FedAvg(FedAvg), and the average performance of decentralized pre-trained local models (Local.avg.).
Figure 4: The disease distribution of the divided 5 subgroups, they are highly heterogeneous.
FedAvg can aggregate semantics from local clients, but the performance would be damaged by data heterogeneity. Fig. 5 shows retrieval accuracies of the strategies we have considered, where the FedAvg pre-trained model achieves higher accuracies than those of decentralized ones with a large margin. However, the centralized pre-trained model acts as an upper bound and there still exists a great space for improvement for FedAvg. Table 10 in the appendix shows the detailed result. Further, we examine if updating the aggregated server model in FedAvg would result in distortion in the learned representation space.
Retraining the server model on local datasets would distort its feature encoders. Fig. 6 demonstrates the effect of distortion by fitting local cross-modal alignments of the corresponding client datasets. We consider a set of models that finished the 25th local iteration and their corresponding aggregated server model. While the average accuracy of local models is less than their aggregated model, the average performance of models re-trained on local clients starting from the aggregated model degrades a lot. Table 11 in the appendix shows the detailed results.
Diverse information is aggregated in shallow layers that mainly extract feature semantics. We re-train the first 4 shallow layers of the aggregated server model on local clients’ datasets for 500 iteration steps. As shown in Fig. 7, the performance of retrained models decreases a lot, which means shallow layers that extract feature semantics are distorted by heterogeneous local datasets. Table 15 in the appendix shows the detailed results.
De-centralized pre-trained deep layers that act as the align module is harmful for extracting diverse semantics. We re-train the first 4 shallow layers of the decentralized pre-trained model on the union of client datasets for 500 steps. Although the performance has successfully increased, there is still a large gap between that of the FedAvg pre-trained baselines and them as shown in Fig. 8. Therefore, the frozen deep layers prevent the model from learning more diverse semantics from the scaled dataset. Table 16 in the appendix shows the detailed results.
Overall, from the empirical findings, we conclude that federated multimodal pre-training is sensitive to the data heterogeneity which results in learning distorted cross-modal alignment. Simply averaging the local model weights does not solve this issue essentially. Thus, we are motivated to propose our FedAID.
Appendix B Experiment Details
B.1 Pre-training setup
Following Wang et al. (2022) we utilize the MIMIC-CXR Bigolin Lanfredi et al. (2022) dataset for multi-modal pre-training. This dataset is widely used in the medical multi-modal learning domain, with $227,835$ image-text pairs from $65,379$ patients. Some related works also have imported additional features to image-text pairs to augment the data. However, we only use the image-text pairs for pre-training to make the results more generalizable. The MIMIC-CXR dataset is open access, it can be obtained through MIMIC-CXR Access.
@@ B.1 Pre-training setup
During the pre-training, local clients only have access to their highly heterogeneous datasets. To construct the heterogeneous client datasets, following Yan et al. (2023) we employ the Latent Dirichlet Allocation (LDA) Blei et al. (2003) to divide the MIMIC-CXR dataset into $5$ partitions based on a selected sensitive attribute. For implementation, we import the corresponding attribute information of given image-text pairs from the Chest-Imagenone dataset Wu et al. (2021) and divide local datasets based on disease category. The disease category is a multi-label binary attribute and is transformed into a multi-class label. That’s because the words in the clinical report are highly related to the disease category as illustrated in Fig 9. We set the heterogeneity degree in the LDA algorithm to be 1 and the subgroups of the MIMIC-CXR that are used as local datasets are summarized in Table 4.
Specifically, we select 5 commonly considered diseases Bannur et al. (2023): ’Edema’,’Pleural Effusion’, ’Consolidation’, ’Pneumothorax’, and ’Pneumonia’. We set the non-NaN value to 1 and then set NaN value to 0 to construct a 5-way binary multi-label. Then we get $2^{5}$-category multi-class label, and we run LDA on it. The distribution of the disease of the client dataset is summarized in Table 4.
Figure 9: Illustration of the strong connection between latent variable and the text modality.
@@
We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 client datasets. Each divided dataset consists of train splits and test splits based on the notation of the MIMIC-CXR. We use $404\times A$ 40 for pre-training and set the batch size to 388. We set the learning rate to $2\times 10^{-5}$, the number of communications to 25. We set the uncertainty radius $\rho=1$, hyper-parameter $\alpha=1,\beta=2$ in the main experiments. For each communication, we randomly sample 50 batches of data from the client datasets.
B.2 Downstream tasks
We conduct Few-shot classification, Medical image segmentation, and Image Retrieval downstream tasks to evaluate the generalization ability of the pre-trained model.
Few-shot classification. To verify the efficiency of the pre-trained model on the general medical image task, we test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) Shih et al. (2019), the task is to predict if the image reflects a pneumonia-positive phenomenon. (2) Covidx Wang et al. (2020), which has three categories that COVID-19, non-COVID pneumonia and normal. We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset of downstream tasks, and evaluate the image classification accuracy. For evaluation, we report the classification accuracy of the models on them.
Medical image segmentation. To investigate the transfer ability of our model on fine-grained tasks, we conduct experiments on medical image segmentation on the RSNA Wang et al. (2020) benchmark. We follow Wang et al. (2022) to convert object detection ground truths of RSNA into masks for semantic segmentation. Similar to Huang et al. (2021a), we construct a U-Net framework using a decoder and our pre-trained image encoder. During the fine-tuning, we freeze the encoder and fine-tune the decoder using 1%, 10% of the training data of the downstream dataset, and then use the Dice score to evaluate the performance of our model.
Image retrieval. To verify if the pre-trained models have learned the semantics and alignment relationships between the image and text contained in the pre-training dataset, we conduct the image retrieval downstream task. Specifically, we utilize the validation splits of the local clients to test the image retrieval performance of the aggregated models. For a text from a given batch of input image-text pairs, we calculate the similarities between this text and images in this batch. Then we sort these similarities and select the top-1 and top-5 images from them. If the corresponding image of the text is contained in the selected set, then the image is retrieved correctly. Here we utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models.
Out-of-Distribution testing benchmark. We adapt the EHRXQA Bae et al. (2023) VQA dataset for evaluation. This dataset is built on the meta-data notation of MIMIC-CXR Bigolin Lanfredi et al. (2022) and ChestImaGenome Wu et al. (2021). We merge the question and answer sentences as the text input and use the corresponding image to construct the image-text pairs for the multi-modal retrieval task. Other settings are the same as the ordinary image retrieval task.
B.3 Backbones and Baselines